# When LLMs Play the Telephone Game: Cumulative Changes and Attractors in Iterated Cultural Transmissions

### v2

@@ 3.4 Attractor strength and position
Human cultural evolution shows that cultural traits sometimes evolve towards attractor states, i.e., content that invites convergence even with different starting points kalish_iterated_2007; miton_universal_2015; miton_motor_2020; buskell_what_2017. Therefore, we were interested in whether transmission chains with LLMs would show similar attractor dynamics, and whether these depend on the model and task used in the chain. The concept of cultural attractor is not consistently formalized in the human cultural evolution literature buskell_what_2017. Here, we defined attractors as the theoretical equilibrium point to which the iterated generation process (defined in Eq. 1) may eventually converge. We mathematically define attractors in terms of two properties of interest: its position (i.e., the location in output generation space the process converges toward) and its strength (i.e., the intensity to which generated outputs are pulled toward it). The strength takes values in $[0,1]$, which allows for a continuous notion of an attractor: rather than being a binary concept that either exists or does not, attractors here lie on a spectrum, covering systems without attraction effects (strength=0) to ideal attractors (strength=1). To compute position and strength, we use the simulated data to fit a linear regression predicting the value of a property at the end of the chain (i.e. after $n_{generations}$ as a function of its value in the initial text (Figure S1).
 [^7] Linear regressions were fit using the SciPy library (https://scipy.org/) released under the BSD Licence
For example for a given text property, we fit:
$\textit{property}(generation=n_{generations})=I+s*\textit{property}_{initial},$ (2)
@@
where I is the estimated intercept and s the estimated slope. This enables us to estimate the final output of a new chain starting from the final output of the previous chain as:
$\textit{property}(generation=2*n_{generations})=I+s*\textit{property}(generation=n_{generations}).$ (3)
The fitted linear regression thus allows to define a recurrent relationship between the output of a chain as a function of the output of the previous chain:
$\textit{property}(generation=k*n_{generations})=I+s*\textit{property}(generation=(k-1)*50).$ (4)
This relationship is a linear recurrence sequence which can be rewritten as:
${property}(generation=k*n_{generations})=s^{k}*(\textit{property}_{initial}-l)+l,$ (5)
@@
where $l=\frac{I}{1-s}$. If $|s|<1$, then the sequence converges, its limit is $l$ and its convergence rate is $1-s$. We can therefore use the estimated relationship Eq. 2 to determine if an attractor exists ($|s|<1$) and, if so, estimate its position $l=\frac{I}{1-s}$ and strength $1-s$.
@@
To validate that these theoretical fixed points correctly capture attraction dynamics, we estimated their positions using only data from the first 10 generations of each chain, and compared the predictions with the actual output after 50 generations. Visual inspection of the results (Appendix B) confirmed that our method is suited for estimating the strength and position of attractors.
@@ 4 Results
Figure 2: Evolution of the distribution of text properties in generated texts across generations for different models and tasks. We here represent the distribution of each of the four properties at each generation, for each model and task. These distributions thus represent the properties observed in the set of 100 transmission chains (20 initial texts * 5 seeds) for each model and task. For each property, task and model, the 50 generations are arranged vertically, with first generations at the top and last generations at the bottom. This representation allows to capture how iterated transmissions shift the distributions toward certain values, and how quickly this shift happens.
For each of the 5 models, 3 tasks, and 20 initial texts, we ran 5 transmission chains with 50 transmission steps. We provide some examples of generated texts in Appendix 1, and complete data on the companion website  [^8] https://sites.google.com/view/llms-play-telephone . By extracting the properties of generated texts at each generation of each chain, we can study the evolution of these properties through generations, measure how they are affected by interactions beyond single-turn effects, as well as detect and characterize theoretical attractors.
@@ 4.1 Qualitative analysis of property evolutions over generations
In Figure 2, we show the evolution of property distributions over generation for each model and task. This reveals important difference depending on the analyzed property, the task and the model. For instance, we observe that toxicity converges very quickly to a very narrow peak centered around 0. This is very different from the evolution of positivity, for which the initial distribution appears to be quite preserved for the Rephrase task (Figure 2, first row), while less constrained tasks such as Take inspiration (second row) and Continue (third row) lead to more visible changes. Interestingly, we observe that for Llama3-8B (blue) and Llama3-70B (pink), the distribution of positivity values converges to a bimodal distribution, while distributions are unimodal for other models. In some cases, we also observe that different models lead the distributions to be shifted in opposite directions. For instance when looking at the evolution of texts length, using GPT3.5 (green) or Llama3-8B (blue) leads text to become on average shorter, while using Mixtral-8x7B shifts the distribution towards greater lengths.
@@ 4.2 To what extent do multi-turn transmissions affect the evolution of properties?
Qualitative analyses from the previous section appear to suggest that multi-turn transmissions lead texts to acquire different properties compared to single-turn settings. To quantitatively evaluate this observation, we use Kolmogorov–Smirnov (KS) tests ks_test to estimate the compare property distributions after a single interaction and after multiple interactions. In Figure 3, we report for each model, task and property the p-value of the KS test for the null hypothesis $H_{0}$: "The text properties at generation $i$ are sampled from the same distribution as the text properties after generation 1". Across most instances, we observe that $H_{0}$ can be confidently rejected, indicating that property distributions become significantly different after multi-turn transmissions compare to single-turn transmissions. We observe that this is more often the case for less constrained tasks (Take Inspiration and Continue, second and third columns) than for more constrained task (Rephrase, first column). This finding confirms that studying single-turn interactions is in general not sufficient for analyzing the properties of interacting LLMs’ outputs. This warrants a more detailed account of the cultural dynamics across iterated interactions among LLMs.
@@ 4.3 What influences the presence, strength, and position of attractors?
Visual inspection of the evolution of text properties as presented on Figure 2 indicate that multi-turn transmissions lead distributions to become skewed toward certain values, which suggests the presence of attractors. The task assigned to a chain and the model type populating it appear to influence the position of those attractors, as well as their strength (i.e. how quickly do shifts in distributions happen). To have quantitative measures of attractors strengths and positions, we use the method described in 3.4 and Figure S1. By fitting a linear regression predicting the value of a given property at the end of the chain as a function of the property in the initial text, we can represent for each property, model and task the position and strength of the attractors, provided they exist. Figure 4 presents the estimated strengths and positions of attractors, and fitted linear regressions are provided as supplementary material in Figure S2.
For all combinations of property, task and model, we found that the recurrent relationship defined by the fitted linear regression converges. This means that all conditions admit a theoretical attractor as defined in Section 3.4. As for the strength and position of these theoretical attractors, Figure 4 already allows to notice some tendencies. For instance, it seems that less constrained tasks (e.g. Continue) lead to stronger attractors than more constrained tasks (e.g. Rephrase). To better disentangle the respective contributions of model type, task and property on attractors, we fitted Bayesian models predicting attractor Strength as a function of Task, Model and Property, and predicting attractor Position as a function of Task and Model, for each of the four considered properties. We report 95% Credible Intervals (CI), and provide details about statistical analyses in Appendix B.
@@
We find a strong effect of Task on attractor strength: Continue leads to significantly stronger attraction than Take Inspiration (95% CI: [0.0511;0.254]) , itself leading to significantly stronger attraction than Rephrase (95% CI: [0.0326; 0.2377]). This confirms our observation that less constrained tasks lead to stronger attraction than more constrained tasks. Different properties are also found to display different sensitivity to attraction effects. We detect that toxicity possesses significantly stronger attractors than positivity (95% CI: [0.1158; 0.3515]), difficulty (95% CI: [0.1072; 0.3417]) and length (95% CI: [0.173; 0.407]). As for the effect of model, we observe significantly weaker attraction for Llama3-70b compared to GPT3.5 (95% CI: [-0.3215; -0.059]) , Llama3-8B (95% CI: [-0.2954; -0.0326]) and Mixtral-8x7b (95% CI: [-0.3163; -0.0541]) .
@@
We next analyze the position of the attractors. We observe that the attractor for toxicity is significantly higher (i.e the attractor state has higher toxicity) for Llama3-8b than for GPT3.5 and Mixtral-8x7b, and significantly higher for Llama3-70b than for GPT3.5, Mistral-7b and Mixtral-8x7b. It is also higher for Continue than for Rephrase. For positivity, we found that the position of the attractor was significantly lower for Llama3-8b than for GPT3.5, Mistral-7b and Mixtral-8x7b, and that the task Take inspiration and Continue both led to significantly higher positivity than the Rephrase task.
@@
Figure 4: Attractors strength and position. The heigth of the bars represent the position (top row) and strength (bottom row) of theoretical attractors estimated using the method described in Section 3.4, for each property (columns), task, and model. Less constrained tasks, such as Continue, appear to produce stronger attractors than more constrained tasks, such as Rephrase. Attractors appear to be stronger for toxicity than for length. Finally, we can notice that the position of attractors appears to vary between models. For instance, the attractor for difficulty apperars to be higher for Llama3-8b and Llama3-70b than for other models. Statistical analyses allowed to quantify these differences and are presented in Section 4.3.
Effect of temperature
@@
Temperature is a hyper-parameter that controls the stochasticity in LLMs outputs. Our experiments indicate that increasing the temperature appears to lead to stronger attractors only in constrained tasks (Rephrase and Take inspiration), but not in the more open- ended task Continue (Figure 5a). One interpretation might be that increasing temperature relaxes constraints on the content that can be produced, thus leading to stronger attractors. This effect would therefore be more significant for tasks that are quite constrained.
Effect of fine-tuning
@@
We observe that fine-tuning can shift the attractors’ positions (Figure 5b): the attractor for length is lower, and the attractor for difficulty higher, for Instruct models compared to Base models. For toxicity and positivity, we do not observe very significant shifts. This may indicate that Base models were already quite aligned with human preferences in terms of toxicity and positivity even before fine-tuning, possibly due to techniques such as data curation. Length and Difficulty may have been less targeted by such techniques, and fine-tuning may therefore have shifted the attractors’ positions towards human preferences. As for the strength of attractors, we observe that fine-tuning seems to increase attraction strength for toxicity, but to reduce it for difficulty. This may suggest that for properties on which humans have strong, uniform preferences (such as toxicity), fine-tuning leads to stronger attractors, while for properties on which humans have weaker and more heterogeneous preferences, it mitigates the strength of attractors that came from the training data. This analysis provides insight about how attractors are formed, and reveals that both the choice of training data and fine-tuning processes may impact LLMs in ways that only become significant in the case of multi-turn behaviors.
@@ 5 Discussion
While current studies analyzing the outputs of LLMs are restricted to a single prompt-output interaction, we borrowed the methodology from studies on human cultural evolution to address how cultural content may evolve over transmission chains with LLMs. This resulted in a series of telephone game experiments assessing the evolution of cultural content in LLMs as a function of models, instructions, and text properties. Our results reveal that several changes in generated content appear after multiple iterations. For example, we observed that the difficulty of a provided text was preserved after an LLM was prompted to elaborate it a single time, but changed dramatically after the text was processed iteratively by a chain of LLMs.
@@
Figure 5: Effect of temperature (a) and fine-tuning (b) on attractors. (a) Attractor positions (top row) and strength (bottow row) for different values on temperature (x-axis), for model Llama3-8B-Instruct. The main visible effect is that increasing temperature increases attraction strength for tasks Rephrase and Take Inspiration, but not for Continue. (b) Attractor positions (top row) and strength (bottow row) for Base and Instruct versions of Mixtral-8x7B and Llama3-70B. The most visible effects are that fine-tuning increases the strength of the attractor for toxicity, increases the position of the attractor for difficulty, and decreases the position of the attractor for length.
Those qualitative observations were confirmed by statistical tests, where we found that multi-turn interactions lead the distributions of text properties to become significantly different distributions obtained after single-turn interactions.
@@
By comparing the properties of input texts to those of texts produced by transmission chains spanning several generations of LLMs, we identified property-specific patterns in the convergence of LLM dynamics toward attractor states. Using this method, we found that some properties (e.g. toxicity) display quicker convergence rates toward attractors, and that the position of these attractors varies between models. Moreover, our results suggest that more open-ended tasks (e.g. Continue) lead to quicker convergence toward attractors than more constrained tasks (e.g. Rephrase). Studying attraction dynamics in systems of LLMs may therefore be particularly relevant to situations in which LLMs are used to simulate artificial societies de_zarza_emergent_2023; park_generative_2023; park_social_2022; xiao_simulating_2023; hua_war_2023; vezhnevets_generative_2023; chuang2023simulating, where they are often granted a relatively high freedom in order to witness emergent behaviors. Evaluating the effect of temperature on attractors supports this view that allowing more room for variation leads to stronger attractors, as we found that increasing temperature increases attraction strength only for quite constrained tasks.
To understand the role of fine-tuning in shaping the strength and position of these attractors, we compared Base models with their Instruct versions. This revealed that fine-tuning can affect both attractors strength (e.g. for toxicity) and position (e.g. for difficulty and length).
@@
Limitations and future work
As the study of the cultural dynamics of generative agents is an emerging research area, our setting involved several simplifications. While we focused on linear transmission chains, real-world interactions typically involve networks of senders and receivers. Human studies have shown that network size fay_increasing_2019; richerson_group_2013; andersson_group_2014; baldini_revisiting_2015; derex_experimental_2013 and structure raviv_role_2020; kirby_cumulative_2022; derex_partial_2016; de_pablo_understanding_2022; derex_cumulative_2020 influence cultural evolution. Following some initial endeavors nisioti_social_2022; perez_cultural_2024, future work may assess similar effects in machine networks. To investigate model- and task-specific biases, we studied transmission chains in homogeneous settings, where agents belonged to the same model type and received the same instructions, but future studies may address cultural dynamics in heterogeneous populations of LLMs prompted with various instructions. While we focused on LLM and, therefore, text outputs, similar studies may be run to address the properties of various generative tools (e.g., for image generation). In the future, researchers may also address hybrid networks in which humans and LLMs interact — a scenario that is becoming increasingly relevant as generative tools become more widespread and which may, in turn, shape the future of human cultural evolution. machine_culture.

### v3

@@ 3.4 Attractor strength and position
Human cultural evolution shows that cultural traits sometimes evolve towards attractor states, i.e., content that invites convergence even with different starting points (Kalish et al., 2007; Miton et al., 2015; 2020; Buskell, 2017). Therefore, we were interested in whether transmission chains with LLMs would show similar attractor dynamics, and whether these depend on the model and task used in the chain. The concept of cultural attractor is not consistently formalized in the human cultural evolution literature (Buskell, 2017). Here, we defined attractors as the theoretical equilibrium point to which the iterated generation process (defined in 3.1) may eventually converge. We mathematically define attractors in terms of two properties of interest: its position (i.e., the location in output generation space the process converges toward) and its strength (i.e., the intensity to which generated outputs are pulled toward it). The strength takes values in $[0,1]$, which allows for a continuous notion of an attractor: rather than being a binary concept that either exists or does not, attractors here lie on a spectrum, covering systems without attraction effects (strength=0) to ideal attractors (strength=1). To compute position and strength, we use the simulated data to fit a linear regression predicting the value of a property at the end of the chain (i.e. after $n_{generations}$ as a function of its value in the initial text (Figure S1).
@@
For example for a given text property, we fit: $\textit{property}(generation=n_{generations})=I+s*\textit{property}_{initial}$, where I is the estimated intercept and s the estimated slope. This enables us to estimate the final output of a new chain starting from the final output of the previous chain as: $\textit{property}(generation=2*n_{generations})=I+s*\textit{property}(generation=n_{generations}).$
@@
The fitted linear regression thus allows to define a recurrent relationship between the output of a chain as a function of the output of the previous chain: $\textit{property}(generation=k*n_{generations})=I+s*\textit{property}(generation=(k-1)*50).$ This relationship is a linear recurrence sequence which can be rewritten as: ${property}(generation=k*n_{generations})=s^{k}*(\textit{property}_{initial}-l)+l$, where $l=\frac{I}{1-s}$. If $|s|<1$, then the sequence converges, its limit is $l$ and its convergence rate is $1-s$. We can therefore use the estimated relationship to determine if an attractor exists ($|s|<1$) and, if so, estimate its position $l=\frac{I}{1-s}$ and strength $1-s$.
@@
To validate that these theoretical fixed points correctly capture attraction dynamics, we estimated their positions using only data from the first 10 generations of each chain, and compared the predictions with the actual output after 50 generations. Visual inspection of the results (Appendix B.5) confirmed that our method is suited for estimating the strength and position of attractors.
@@ 4 Results
Figure 2: Evolution of the distribution of text properties across generations. We here represent the distribution of each of the four properties at each generation, for each model and task. These distributions thus represent the properties observed in the set of 100 transmission chains (20 initial texts * 5 seeds) for each model and task. For each property, task and model, the 50 generations are arranged vertically, with first generations at the top and last generations at the bottom.
For each of the 6 models, 3 tasks, and 20 initial texts, we ran 5 transmission chains with 50 transmission steps. In the main experiment, each chain is composed of a population of agents sharing the same underlying model. In an additional experiment, we also studied chains where different models interact with one another (Appendix B.2.4). We provide some examples of generated texts in Appendix Section 1, and complete data on the companion website  [^4] https://sites.google.com/view/telephone-game-llm. By extracting the properties of generated texts at each generation of each chain, we can study the evolution of these properties through generations, measure how they are affected by interactions beyond single-turn effects, as well as detect and characterize theoretical attractors.
@@ 4.1 Qualitative analysis of property evolutions over generations
In Figure 2, we show the evolution of property distributions over generation for each model and task. This reveals important difference depending on the analyzed property, the task and the model. For instance, we observe that toxicity converges very quickly to a very narrow peak centered around 0. This is very different from the evolution of positivity, for which the initial distribution appears to be quite preserved for the Rephrase task (Figure 2, first row), while less constrained tasks such as Take inspiration (second row) and Continue (third row) lead to more visible changes. Interestingly, we observe that for Llama3-8B and Llama3-70B, the distribution of positivity values converges to a bimodal distribution, while distributions are unimodal for other models. In some cases, we also observe that different models lead the distributions to be shifted in opposite directions. For instance when looking at the evolution of texts length, using GPT3.5 or Llama3-8B leads text to become on average shorter, while using Mixtral-8x7B or GPT-4o-mini shifts the distribution towards greater lengths.
@@ 4.2 To what extent do multi-turn transmissions affect the evolution of properties?
Qualitative analyses from the previous section appear to suggest that multi-turn transmissions lead texts to acquire different properties compared to single-turn settings. To quantitatively evaluate this observation, we use Kolmogorov–Smirnov (KS) tests (Massey, 1951) to estimate the compare property distributions after a single interaction and after multiple interactions (see Section B.3.1 for justification of why this method is adequate). In Figure 3, we report for each model, task and property the p-value of the KS test for the null hypothesis $H_{0}$: "The text properties at generation $i$ are sampled from the same distribution as the text properties after generation 1". Across most instances, we observe that the p-values steadily decrease, indicating that observing the given distribution under the null hypothesis becomes less and less likely with generations. We observe that this is more often the case for less constrained tasks (Take Inspiration and Continue, second and third columns) than for more constrained task (Rephrase, first column). This finding confirms that studying single-turn interactions is in general not sufficient for analyzing the properties of interacting LLMs’ outputs. This warrants a more detailed account of the cultural dynamics across iterated interactions among LLMs.
@@ 4.3 What influences the presence, strength, and position of attractors?
Effect of model, task and property Visual inspection of the evolution of text properties as presented on Figure 2 indicate that multi-turn transmissions lead distributions to become skewed toward certain values, which suggests the presence of attractors. The task assigned to a chain and the model type populating it appear to influence the position of those attractors, as well as their strength (i.e. how quickly do shifts in distributions happen). To have quantitative measures of attractors strengths and positions, we use the method described in 3.4 and Figure S1. Figure 4 presents the estimated strengths and positions of attractors, and fitted linear regressions are provided as supplementary material in Figure S2. For all combinations of property, task and model, we found that the recurrent relationship defined by the fitted linear regression converges. This means that all conditions admit a theoretical attractor as defined in Section 3.4. To better disentangle the respective contributions of model type, task and property on attractors position and strength, we fitted bayesian models predicting attractor Strength as a function of Task, Model and Property, and predicting attractor Position as a function of Task and Model, for each of the four considered properties. We use 95% Credible Intervals (CI) to assess significance. Those confidence intervals,as well as details of statistical models, are provided in Appendix section B.
@@
We find a strong effect of Task on attractor strength: Continue leads to significantly stronger attraction than Take Inspiration , itself leading to significantly stronger attraction than Rephrase . This supports our observation that less constrained tasks lead to stronger attraction than more constrained tasks. To specifically test this hypothesis, we conducted additional experiments manipulating only the room for variation allowed in the task, which confirmed this hypothesis (Figure S5).
Different properties are also found to display different sensitivity to attraction effects. We detect that toxicity possesses significantly stronger attractors than positivity, difficulty and length. As for the effect of model, we observe significantly weaker attraction for Llama3-70b compared to GPT3.5, Llama3-8B and Mixtral-8x7b. GPT-4o-mini also displayed singificantly weaker attractors than GPT3.5, Llama3-8B and Mixtral-8x7b.
@@
As for the position of the attractors, we found that the position of the attractor for positivity was significantly lower for Llama3-8b than for GPT3.5, GPT-4o-mini, Mistral-7b and Mixtral-8x7b, and that the task Take inspiration and Continue both led to significantly higher positivity than the Rephrase task.
@@
Figure 4: Attractors strength and position. The heigth of the bars represent the position (top row) and strength (bottom row) of theoretical attractors estimated using the method described in Section 3.4, for each property (columns), task, and model. Less constrained tasks, such as Continue, appear to produce stronger attractors than more constrained tasks, such as Rephrase. Attractors appear to be stronger for toxicity than for length. Finally, we can notice that the position of attractors appears to vary between models.
Robustness checks To verify the robustness of our results, we verified that they hold when using a larger set of initial texts (Figure S3) and when using different phrasings of the instructions prompts (Figure S4). We observed similar trends and values as in the main experiment, which confirms the robustness of our results.
@@
Effect of temperature Increasing the temperature appears to lead to stronger attractors only in constrained tasks (Rephrase and Take inspiration), but not in the more open- ended task Continue (Figure 5a). One interpretation might be that increasing temperature relaxes constraints on the content that can be produced, thus leading to stronger attractors. This effect would therefore be more significant for tasks that are quite constrained.
@@
Effect of fine-tuning We observe that fine-tuning can shift the attractors’ positions (Figure 5b): the attractor for length is lower, and the attractor for difficulty higher, for Instruct models compared to Base models. For toxicity and positivity, we do not observe very significant shifts. This may indicate that Base models were already quite aligned with human preferences in terms of toxicity and positivity even before fine-tuning, possibly due to techniques such as data curation. Length and difficulty may have been less targeted by such techniques, and fine-tuning may therefore have shifted the attractors’ positions towards human preferences. As for the strength of attractors, we observe that fine-tuning seems to increase attraction strength for toxicity, but to reduce it for difficulty. This may suggest that for properties on which humans have strong, uniform preferences (such as toxicity), fine-tuning leads to stronger attractors, while for properties on which humans have weaker and more heterogeneous preferences, it mitigates the strength of attractors that came from the training data. This analysis provides insight about how attractors are formed, and reveals that both the choice of training data and fine-tuning processes may impact LLMs in ways that only become significant in the case of multi-turn behaviors.
@@ 5 Discussion
While current studies analyzing the outputs of LLMs are restricted to a single prompt-output interaction, we borrowed the methodology from studies on human cultural evolution to address how cultural content may evolve over transmission chains with LLMs. This resulted in a series of telephone game experiments assessing the evolution of cultural content in LLMs as a function of models, instructions, and text properties. Our results reveal that several changes in generated content appear after multiple iterations. For example, we observed that the difficulty of a provided text was preserved after an LLM was prompted to elaborate it a single time, but changed dramatically after the text was processed iteratively by a chain of LLMs. Those qualitative observations were confirmed by statistical tests, where we found that multi-turn interactions lead the distributions of text properties to become significantly different distributions obtained after single-turn interactions.
@@
Figure 5: Effect of temperature (a) and fine-tuning (b) on attractors. (a) Attractor positions (top row) and strength (bottow row) for different values on temperature (x-axis), for model Llama3-8B-Instruct. The main visible effect is that increasing temperature increases attraction strength for tasks Rephrase and Take Inspiration, but not for Continue. (b) Attractor positions (top row) and strength (bottow row) for Base and Instruct versions of Mixtral-8x7B and Llama3-70B. Fine-tuning appears to increase the strength of attraction for toxicity, increases the position of the attractor for difficulty, and decreases the position of the attractor for length.
@@
By comparing the properties of input texts to those of texts produced by transmission chains spanning several generations of LLMs, we identified property-specific patterns in the convergence of LLM dynamics toward attractor states. Using this method, we found that some properties (e.g. toxicity) display quicker convergence rates toward attractors, and that the position of these attractors varies between models. Moreover, our results reveal that more open-ended tasks (e.g. Continue) lead to quicker convergence toward attractors than more constrained tasks (e.g. Rephrase). Studying attraction dynamics in systems of LLMs may therefore be particularly relevant to situations in which LLMs are used to simulate artificial societies (de Zarzà et al., 2023; Park et al., 2023; 2022; Xiao et al., 2023; Hua et al., 2023; Vezhnevets et al., 2023; Chuang et al., 2023), where they are often granted a relatively high freedom. Evaluating the effect of temperature on attractors supports this view that allowing more room for variation leads to stronger attractors, as we found that increasing temperature increases attraction strength only for quite constrained tasks. To understand the role of fine-tuning in shaping the strength and position of these attractors, we compared Base models with their Instruct versions. This revealed that fine-tuning can affect both attractors strength (e.g. for toxicity) and position (e.g. for difficulty and length).
@@
Limitations and future work As our results indicate that both training data and fine-tuning affect attractors properties, future work should further explore how different fine-tuning and data curation methods shape multi-turn dynamics. While we focused on linear transmission chains, real-world interactions typically involve networks of senders and receivers. Human studies have shown that network size (Fay et al., 2019; Richerson, 2013; Andersson and Read, 2014; Baldini, 2015; Derex et al., 2013) and structure (Raviv et al., 2020; Kirby and Tamariz, 2022; Derex and Boyd, 2016; de Pablo et al., 2022; Derex and Mesoudi, 2020) influence cultural evolution. Following some initial endeavors (Nisioti et al., 2022; Perez et al., 2024; Lai et al., 2024), future work may assess similar effects in machine networks, and in particular investigate how attraction dynamics are shaped by network-based strategies such as dynamic network structures (Nisioti et al., 2022) or free-formed decentralized networks (Lai et al., 2024). To investigate model- and task-specific biases, we focused on homogeneous transmission chains where agents belonged to the same model type and received the same instructions. Future studies may address cultural dynamics in heterogeneous populations of LLMs prompted with diverse instructions. Our study could also be extended to hybrid networks in which humans and LLMs interact — a scenario that is becoming increasingly relevant as generative tools become more widespread and which may, in turn, shape the future of human cultural evolution (Brinkmann et al., 2023).
Reproducibility Statement
The code for reproducing simulations, statistical analyses, and for generating all figures can be found on our GitHub repository: https://github.com/jeremyperez2/TelephoneGameLLM. This repository also contains the data generated when running the experiments. This data can also be accessed from the companion website (https://sites.google.com/view/telephone-game-llm), using the Data Explorer tool.
Ethics Statement
The results presented in this paper highlight that multi-turn interactions can give rise to potential harmful biases that cannot be detected by only studying the outcome of single-turn interactions. By introducing novel conceptual and methodological tools, we provide ways of quantifying the properties of multi-turn interaction systems with respect to the extent to which they magnify or reduce these biases. It is then the responsibility of end-users to make an ethical use of these tools, for instance by using them to ensure that LLMs remain aligned with ethical standards even after repeated interactions.
Acknowledgements
This research was partially funded by the French National Research Agency (ANR, project ECOCURL, Grant ANR-20-CE23-0006). This work benefited from access to the Jean Zay (Idris) supercomputer associated with the Genci grant A0151011996. We also thank Chris Foulon, Marcela Ovando-Tellez and Joan Dussauld who participated in the hackathon Hack1Robo during which this project originated. M.D. acknowledges Institute for Advanced Study in Toulouse (IAST) funding from the French National Research Agency (ANR) under grant no. ANR-17-EURE-0010 (Investissements d’Avenir programme).